OpenStudy (anonymous):

write an equation in slope-intercept form for the line that satisfies the condition: x-intercept=5, y-intercept=3

OpenStudy (anonymous):

slope intercept is in the form of: y = mx + b to write an equation in the form of slope intercept, you leave x and y as x and y, and substitute in the slope for m and the y intercept for b

OpenStudy (anonymous):

You already know that the y intercept is 3. That gives you your value for b, 3. The intercepts give you more information. From the intercepts you know that two points on the line are (5, 0) and (0, 3). Slope is change in y divided by change in x (also sometimes called rise over run). You can use (y1 - y2) / (x1-x2) or (y2 - y1) / (x2 - x1). In this case the slope is (0 - 3) / (5 - 0) = -3/5. That is your value for m.

OpenStudy (anonymous):

Using those, your equation in slope intercept form is y = -3/5x + 3

OpenStudy (anonymous):

Thanks so much!